find the distance between each pair of parallel lines
x+3y=-14
1x+3y=-14
-3y -14
----------
1x =-14
-- ---
1 1
---------
x =-14
x+3y=6
1x+3y=6
-3 -3
--------
1x = 3
-- -
1 1
--------
x = 3

Answer by sophmore(3)

x+3y=6
1x+3y=6
-3 -3
--------
1x = 3
-- -
1 1
--------
x = 3

x+3y=-14
1x+3y=-14
-3y -14
----------
1x =-14
-- ---
1 1
----------
x =-14